OpenStudy (anonymous):

A number decreased by 32 is -58. Find the number. The number is ____. (Give only the value of the number as your answer.)

OpenStudy (anonymous):

-26

OpenStudy (anonymous):

if you have a number (= X ) and you decrease 32 so you can get -58, you just have to add 32 to -58 to find your number

OpenStudy (ganpat):

x - 32 = -58 x = -58 + 32 = -26...

OpenStudy (anonymous):

Thank You! Can you guys continue to help me

OpenStudy (anonymous):

x=-26